OpenStudy (anonymous):

What is the quotient when x^3 -5x+3x - 8 is divided by x-3? A.x^2-8x+27 B. x^2-2x-3 C.x^2-8x+27 - 89/(x-3) D.x^2 -2x - 3 - 17/(x-3)

OpenStudy (dumbcow):

have you learned synthetic division?

OpenStudy (dumbcow):

http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (dumbcow):

3 | 1 -5 3 -8 3 -6 -9 -------------- 1 -2 -3 -17

OpenStudy (anonymous):

yes I just finished the lesson and I can't seem to figure out this problem. I know you have to put them in decending order, then multiply and divide. but my answer is nothing close to the choices they give.

OpenStudy (anonymous):

So is it D?

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

So what about 2x^3+x+3 divided by x+1

OpenStudy (dumbcow):

-1 | 2 0 1 3 ------------ you finish it

OpenStudy (anonymous):

what about What is f(x)/g(x) when f(x)= -2x^3+x^2+16-15 and g(x) = x+3?